# How Exclusive are Ethereum Transactions?

### v1

@@ I Can Haz Block? 31 1 / 24 0 / 0 31 1 / 24 0 / 0
We next examine how exclusive transactions from the same sender differ when sent to multiple builders. Specifically, we identify transactions originating from the same to address but submitted to different builders within the same auction cycle. [^12] As noted, multiple from addresses may use the same to address. Defining a sender by its to address, therefore, makes it more likely that we detect similarities among transactions from the same sender but with different hashes. In this sense, using to addresses provides a stricter test of our hash-based classification of transactions as exclusive, private, or public. We then re-execute these transactions to extract their logs—the sequence of contract calls and outputs each transaction generates. In total, we identify 665 such transactions producing 279 unique logs. Nearly all of them (630 transactions) are swaps in DEX pools, primarily on Uniswap V2, Uniswap V3, Sushiswap, Pancakeswap, or Curve.
We categorize these transactions based on their logs and fees:
We find 171 partially exclusive transactions (59 unique logs) with a degree of exclusivity equal to zero—that is, the sender submitted the same transaction (same log, same fee) to at least two builders. We found another 494 partially exclusive transactions (220 unique logs) with different fees across builders. Among these, the average degree of exclusivity is 0.004178 ETH (or 23% of the value of the second-highest fee).
Finally, we identify 1,179 “similar” exclusive transactions grouped into 162 log sets (861 unique logs), where transactions differ only in the swap amount. Comparing the fees attached to these transactions, we find that the average gap between the highest payment to one builder and the second-highest payment to another builder is almost 70% of the second-highest payment. This suggests that the differences between these transactions are economically meaningful.
The takeaway is that not all transactions that appear “exclusive” at the hash level remain exclusive once we examine their logs: some are duplicates submitted to multiple builders, while others are variations of the same strategy. Quantitatively, however, these cases are limited—only 665 partially exclusive transactions out of 5,576 in total. Also, if we change the classification of partially exclusive and similar transactions to “private”, we find that 77.2% of fees paid by transactions that are included on-chain are from exclusive transactions (vs 84% with our initial classification). The main result, therefore, stands: exclusive transactions remain the key driver of builders’ profitability.
@@ 5 Conclusion
This paper provides the first systematic analysis of non-winning Ethereum blocks, allowing us to distinguish between public, private, and exclusive transactions. We show that, although exclusive transactions represent a minority of all transactions, they account for the vast majority of builder revenues—roughly 84% of fees in winning blocks. Exclusive relationships between senders and builders are widespread but account for a minimal share (approximately 7%) of the value of exclusive transactions included on-chain. These findings help explain why block-building is so concentrated, with a small number of builders dominating the market.
We also find that a significant share of transactions that appear exclusive in one bidding cycle become private or even public in later cycles. Examining transaction logs reveals that some “exclusive” transactions are actually duplicates or variations of the same strategy sent to multiple builders. Yet these cases are quantitatively limited: exclusivity measured at the hash level still provides an accurate picture of the determinants of builders’ profitability.
Our results contribute to the debate on the extreme concentration of Ethereum’s builder market. By showing that exclusive transactions generate the vast majority of builder revenues, we provide support to the view that exclusivity is an important driver of concentration and helps explain the dominance of a few builders.

### v2

@@ I Can Haz Block? 31 1 / 24 0 / 0 31 1 / 24 0 / 0
4.3 Transactions’ transitions.
Among the 1,330 transactions observed across multiple bidding cycles, public transactions remain public in all subsequent cycles. Only a handful of private transactions (13 out of 208) later appear in the public mempool. By contrast, the majority of exclusive transactions transition to private or even public. Of the transactions that begin as exclusive, 162 remain exclusive across all observed cycles, while 174 transition to private. [^12] We also find 33 transactions that were initially exclusive to a builder and later included in a subsequent cycle won by the same builder. Because those cycles are not in our primary dataset, we cannot determine whether the transactions remained exclusive or became private. In addition, 12 transactions that were initially exclusive eventually surfaced in the public mempool. Nine of these were exclusive for two cycles and three for three cycles. In four cases, after appearing in the public mempool, these transactions were added on chain by the same builder that had them as exclusive to start with (Titan). All appear to be user-generated: three used the 0x router and four the 1inch router. The existence of these transactions is problematic, as they were first delayed and later exposed in the public mempool, leaving users vulnerable to attacks (and one of them appears to have been sandwiched). [^13] These transactions are: 0x1daea1584d684385fb25209bad3a49c54a1e2c500e27543334c7b3e695a12ffb, 0x6605c1c36fcd3d696844519d6ff7199fe4ddd34fb4f3e0ed4da71aedfcf1c633, 0x3dadbab8e85b6e48a2c33f9029ea4ebebdd3def8b9d907faff15ef81527de968, 0x9c7364eddb94fe659bc8104585ff86a8a234480fe00a9f9aa3eb410e572ebdf2, 0x9d476dbc524651cbe26bcafab99f6fe40a823a5588fa5cd5002e5b414a8a2f8c, 0x417440b083bdcb529663c6615c32d89a92beb7863f1a5d6b12bca5b94f4a0f87 (sandwiched according to https://www.zeromev.org/block?num=21322643), 0xe56caee27e79dee8a4c7ed6f945ccbc2cc733d761e19bd68329dc364c61e73a7, 0xae5b88bc0bffa04811eda0dfdcd7224599e2757339e18b2d57f4f40d61902a4c, 0xd41139b67a193045b4d60266c084f6dadd0cb95dccd7b2a06ea68653d6a313ec, 0x597fc0cb0610d1e3ea2baf0f6f90d0d741a6429bfdd9b375bfd2493b70d41dcd, 0x44e8105d5c672195877dc1ed970511e0f0a15fc25cbcafe71d51d428d54c215d, 0xd74718e6d8d58bc33334c0c2f6642d9dec49496329329d8fdcb1e698e81abc94.
Of Titan’s 17 cases, 8 came through the Flashbots private mempool. Five originated from a single searcher address1414140x89a99a0a17d37419f99cf8dc5ffa578f3cdb58b5 and appear in Flashbots Protect’s dataset as shared with more than 15 builders. These transactions remained exclusive to Flashbots Builder for about 2.6 seconds before entering Titan’s blocks. Another five transactions from the same address follow the same pattern (initial exclusivity to Flashbots, later appearance in Titan’s blocks) but were never included on-chain. It is plausible that they were also shared via Flashbots Protect, which records only transactions that are eventually included.
In summary, transaction sharing between builders during the bidding cycle does occur but is relatively rare in our data. When it does occur, it involves transactions submitted via Flashbots Protect that are initially exclusive to the Flashbots builder. We speculate that Flashbots may share such transactions with other builders once it becomes apparent that its own builder is unlikely to win the auction.
@@ 5 Conclusion
Together with its companion paper Canidio and Pahari (2025), this study provides the first systematic analysis of non-winning Ethereum blocks, enabling us to distinguish between public, private, and exclusive transactions. We show that although exclusive transactions represent a minority of all transactions included on chain, they account for the vast majority of builders’ revenues—between 77.2% and 84% of total fees in winning blocks, depending on the methodology. While exclusive relationships between senders and builders are common, they explain only a small share (approximately 7%) of the total value of exclusive transactions included on-chain. By demonstrating that exclusive transactions are responsible for the vast majority builder revenues, our results support the view that exclusivity is a key driver of market concentration and helps explain the dominance of a few builders.
We also find that transactions initially classified as exclusive may later change status to private or even public. Specifically, some transactions submitted via Flashbots Protect are exclusive to the Flashbots builder during the early part of a bidding cycle but are later shared with other builders, thereby transitioning to private. In addition, we identify 12 transactions that remain exclusive to a single builder for several bidding cycles before eventually appearing in the public mempool. These transactions are delayed and then exposed to potential attacks.